OpenStudy (anonymous):

suppose the graph of f contains the point (5,2). Find a point that must be on the graph of g(x)=3+f(7x)

OpenStudy (anonymous):

if 7x=5, we know that f is 2. Then putting this into g you get that x=5/7 and g=5

OpenStudy (anonymous):

How did you get g=5

OpenStudy (anonymous):

I used the fact that when we put x=5/7 the f that is being used in g assume value 2.